OpenStudy (mikezack123):

can you check my work. i just need to make sure im right, medal will be given.

OpenStudy (mikezack123):

y - 3 = 2(x - 6)

OpenStudy (mikezack123):

y=1/2x

OpenStudy (marissalovescats):

What step did you do first? :)

OpenStudy (mikezack123):

@marissalovescats i did a lot of steps but i didnt type them here but i wrote my result above your answer

OpenStudy (marissalovescats):

Your answer is not quite correct.

OpenStudy (marissalovescats):

So walk me through how you solved it please. What is the 1st step to solving this?

OpenStudy (mikezack123):

oops i copied down the equation wrong hold on a sec.

OpenStudy (marissalovescats):

Hahah okay

OpenStudy (mikezack123):

y + 3 =1/2(x + 6)

OpenStudy (mikezack123):

lol my result is y=1/2x

OpenStudy (marissalovescats):

You are correct! Great job!

OpenStudy (mikezack123):

thanks! @marissalovescats :)